i need help to solve the commutative or distrbutive to find equivalent expression for the following:
(x+y)x + (x+y)y

Answer by stanbon(75887)

(x+y)x + (x+y)y
---
Using the distributive law you get:
= x^2+xy + xy+y^2
Combine "like" terms:
= x^2+2xy+y^2
Factor
= (x+y)(x+y)
---------------------------
OR
(x+y)x+(x+y)y
The common factor is (x+y):
= (x+y)(x+y)
====================
Cheers,
Stan H.